# Do Androids Know They’re Only Dreaming of Electric Sheep?

### v1

@@ 3.3 Annotation
We ask annotators to label the minimal hallucinatory spans in each response, if any. These are the smallest text spans that would need to be modified to transform an ungrounded response into a grounded one.
A team of 17 annotators completed a pilot annotation exercise of 50 annotations per task. These pilot annotations were checked and verified by the authors to pinpoint common sources of error and disagreement. This feedback was integrated into a revised set of annotation guidelines, shown in Appendix˜A, and the full dataset was then reannotated by the same team. Each example was annotated once; example annotations are shown in Table˜1. The number of annotated examples is shown in Table˜3. Note that for each dataset, over one-fifth of all annotated responses were annotated as containing at least one hallucinated span (columns H-o and H-s).
@@ 4 Probing
Figure 2: General probe structure. Response tokens that were annotated as hallucinatory are shown in red. For each internal representation layer, and each type of representation (attention output and feed-forward output), a classifier is separately trained on the attention head out-projection and feed-forward hidden states for all representation layers (or just the final token) of a response from a prompted language model. Depending on the probe, these classifiers are then ensembled to make a final prediction of whether each individual token was, or otherwise if the response contained, a hallucination.
Probes Alain and Bengio (2017) are tools used to analyze a neural network’s internal representations. Often taking the form of linear classifiers, they are applied to the internal representations of a network and are trained to discriminate between different types of inputs or outputs. Here, we are interested in detecting hallucinatory outputs during an LLM’s generation steps when conditioned on a prompt that requires a response grounded n the prompt.
@@ 4.1 Design
Let $u$ be a prompt as described at the start of Section 3. To determine whether a response $x$ to a prompt $u$ contains hallucinations, we will probe the hidden states of the transformer language model (i.e., decoder) as it generates $x$. Note that these hidden states depend on $u$ and on the prefix of $x$ that has been generated so far and that they may be reconstructed from the string pair $(u,x)$ via forced decoding. A single transformer layer actually consists of two sublayers: the first transforms a token’s embedding by adding a vector predicted by attention, and the second transforms it again by adding a vector predicted by a feed-forward network. We train a separate probe for each of the $2L$ sublayers, where $L$ is the number of layers.
@@
This probe is a linear classifier using a single hidden state vector (output by some sublayer) as input. Let $x_{i}$ be the $i^{\textrm{th}}$ response token and let $\mathbf{h}_{i}\in\mathbb{R}^{n}$ be the corresponding LLM hidden state at a fixed layer and component. We write $y_{i}\in\{0,1\}$ to indicate whether $x_{i}$ falls within a minimum hallucination span. The linear probe predicts $y_{i}$ by $p(y_{i}=1\mid u,x_{\leq i})=\sigma(\mathbf{w}^{\intercal}\mathbf{h}_{i}+b)$, where $\mathbf{w}\in\mathbb{R}^{n}$ are learned weights, $b\in\mathbb{R}$ is a learned bias, and $\sigma$ is the sigmoid function.
@@
Our final approach is a single feed-forward layer that combines the predictions from our $2L$ separately trained probes (either the linear probes or the attention-pooling probes):
$p(y_{i}\mid u,x_{\leq i})$ $=\sigma\left(\sum_{\mathclap{\begin{subarray}{c}\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;l\in\{1,\ldots,L\}\\ \;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;\;m\in\{\textrm{attention},\textrm{feed-forward}\}\end{subarray}}}\beta_{l,m}\cdot p_{l,m}(y_{i}\mid u,x_{\leq i})\right)$
where $\beta_{l,m}$ are learned weights. The individual probes are frozen before $\beta_{l,m}$ parameters are fit.
@@ 4.2 Training Objectives
To train a probe to predict token-level hallucination indicators $y_{i}$, we minimize negative log-likelihood on the annotations in our dataset $\mathcal{D}$:
$\mathcal{L}(\theta)\propto\sum_{(u,x,y)\in\mathcal{D}}\sum_{i=1}^{|x|}-\log p(y_{i}\mid u,x_{\leq i};\theta)$
@@
We also try training a classifier to predict whether $x$ contains any hallucinations at all; denote this indicator as $y=\bigvee_{i=1}^{|x|}y_{i}\in\{0,1\}$. For this, we simply use an attention-pooling probe over the entire response $x$. To fit the probe parameters, $\theta$, we again minimize negative log-likelihood:
$\mathcal{L}(\theta)\propto\sum_{(u,x,y)\in\mathcal{D}}-\log p(y\mid u,x_{\leq|x|};\theta)$
@@ 4.3 Metrics
Token-level Classification.
For token-level classifiers, it is possible to evaluate the F1 score of their per-token predictions; however, this would give more weight to annotated or predicted spans that are longer in length. Here, we are more interested in span-level F1: specifically, whether we predicted all of the annotated hallucinations (recall), and whether all of the predicted spans were annotated as actual hallucinations (precision). As requiring an exact match of the spans is too severe—Appendix˜B shows how our human annotators often disagreed on the exact boundaries of a hallucinatory span—we modify span-level precision and recall metrics to give partial credit, as follows.
Define a span to be a set of tokens, where tokens from different responses are considered distinct. Let $A$ be the set of all annotated spans in the reference corpus, and let $P$ be the set of all predicted spans on the same set of examples. We compute the recall $r$ as the average coverage of each reference span, i.e., the fraction of its tokens that are in some predicted span. Conversely, we compute the precision $p$ as the average coverage of each predicted span:
$r$ $=\frac{1}{|A|}\sum_{s\in A}\frac{\left|s\cap\left(\bigcup_{\hat{s}\in P}\hat{s}\right)\right|}{|s|}$
$p$ $=\frac{1}{|P|}\sum_{\hat{s}\in P}\frac{\left|\hat{s}\cap\left(\bigcup_{s\in A}s\right)\right|}{|\hat{s}|}$
where $\cup$, $\cap$, and $|\cdot|$ denote set union, intersection, and cardinality respectively. We define F1-Sp (standing for "span, partial credit") as the harmonic mean of $p$ and $r$. Note that F1-Sp micro-averages over examples, ignoring the boundaries between them. Thus, in contrast to F1-R, examples with more spans have more influence on the final F1-Sp metric achieved.
@@ 5.1 Probe Hyperparameters and Training
Single-layer hidden state probes, described in Section 4, are trained under float32 precision with unregularized log loss. A batch size of 100 and Adam optimization Kingma and Ba (2014) with a learning rate of 0.001 and $\beta_{1},\beta_{2}=(0.9,0.999)$ were chosen after a small grid search on hyperparameters. With 70/10/20% train, validation, and test splits, each classifier was trained using early stopping with 10-epoch patience. With the predictions of these single-layer probes as input, the ensemble classifier was trained with the same hyperparameters.
@@ 5.2 Response-Level Baselines
We compare the performance of our probes against a set of contemporary methods used in faithfulness evaluation, measuring test set hallucination classification F1. Specifically: (1) the model uncertainty-based metric of length-normalized sequence log-probability Seq-Logprob Guerreiro et al. (2023); (2) off-the-shelf synthetically-trained sentence-level claim classification FactCC Kryscinski et al. (2020); (3) linguistic feature-based dependency arc entailment DAE Goyal and Durrett (2020); (4) QA-based metrics FEQA Durmus et al. (2020); (5) QuestEval Scialom et al. (2021); (6) the NLI-based metric SummaC Laban et al. (2022); and ChatGPT classifiers Luo et al. (2023) (7) without and (8) with chain-of-thought prompting, with prompts shown in §˜C.1.
@@
OC 0.38 0.78 0.72
Seq-Logprob 0.44 0.81 0.75
FactCC 0.43 - -
DAE 0.37 - -
FEQA 0.35 - -
QuestEval 0.41 - -
SummaCZS 0.46 - -
SummaCConv 0.42 - -
ChatGPT 0.34 0.62 0.37
ChatGPT CoT 0.36 0.68 0.45
PoolingSL 0.43 0.86 0.86
PoolingE 0.45 0.88 0.87
PoolingSL-SYNTH 0.46 0.64 0.66
PoolingE-SYNTH 0.48 0.65 0.69
Human 0.71 0.87 0.87
Table 4: Response-level probe and baseline organic hallucination detection performance across tasks as measured via F1 scores (F1-R) achieved on the test set. SL denotes the best single-layer classifier (chosen on the validation set), E denotes the ensembled classifier, and SYNTH indicates that the classifier was trained on synthetic data and tested on organic hallucinations. The Human row measures the performance of a human annotator who is different from the reference annotator; this was measured on a subset of 50 examples that were triply annotated, averaging over all 6 (annotator, reference annotator) pairs.
Metrics (2)–(6) were not originally intended to detect hallucinations in individual responses, but rather to evaluate a collection of responses for the sake of system comparisons. To adapt to our setting, however, we threshold each of these continuous metrics to classify the individual responses, setting the threshold to maximize validation set F1, following Laban et al. (2022). For sentence-level optimized baselines such as FactCC, response-level evaluation is performed by logical OR-ing sentence-level baseline predictions (obtained after thresholding the raw scores). For CDM, the synthetic training dataset (from BUMP) is smaller than the organic training dataset, so the synth results are at a disadvantage. For CF and E2E, however, we generated a synthetic training dataset that was matched in size.
Refer to captionRefer to captionRefer to caption
Figure 3: Grounding behavior saliency—as measured by probe response-level F1 on test set—of single layer probe classifiers trained on the attention-head out projections and feed-forward states of llama-2-13b, stratified across tasks and layers.
@@
OC 0.09 0.58 0.15
LinearSL 0.24 0.65 0.46
LinearE 0.27 0.66 0.47
PoolingSL 0.32 0.68 0.46
PoolingE 0.34 0.68 0.47
Human 0.61 0.69 0.58
Table 5: Span-level probe hallucination detection performance, as measured via span-level partial match scores (F1-Sp). The Human row is computed as in Table˜4 and is relatively stable across tasks, unlike the systems.
5.3 Results
Response-level Results.
On two datasets, we were actually able to match the interannotator agreement rate (the “Human” row). On these datasets, the probes trained on organic hallucinations worked best (Table˜4). The exception was the CDM dataset, where the organic probes lost slightly to the synthetic probes, which in turn fell well short of human performance. On all datasets, we were able to outperform all baseline systems, showing that probing is a feasible (and efficient) alternative to evaluating LLM outputs when model states are available.
Relative to other baselines that also have some access to model internals—such as Seq-Logprob—we hypothesize that probing-based approaches perform better here due to being able to capture information about model uncertainty that is lost when reduced to the output token probability.
Span-Level Results.
These results appear in Table˜5. Again, the “Human” row serves as a ceiling on how well we can expect our automated system to perform on these datasets, given that even humans do not fully agree. We observe that our ensemble attention-pooling probe nearly meets this ceiling for Conv-FEVER but lags behind for CDM and E2E. Still, even this level of performance is potentially useful for avoiding hallucinations, as we discuss in §˜6.
Layers.
Evaluating the performance of trained single-layer probes (Figure˜3), we see that signals of hallucination are strongest in model middle layers during decoding. Probe performance shows a consistent and gradual increase from the zeroth layer up until layer 10 to 20, after which there is a consistent decline in hallucination classification performance for the subsequent layers.
Response-level Classification, F1-R
Base Model Size CDM CF E2E
7B 0.45 0.88 0.87
13B 0.45 0.88 0.87
70B 0.45 0.89 0.87
Span-level Classification, F1-Sp
Base Model Size CDM CF E2E
7B 0.65 0.77 0.46
13B 0.64 0.77 0.46
70B 0.64 0.76 0.46
Table 6: Response and span-level F1 hallucination detection performance, as measured by F1-R and F1-Sp, respectively, of the PoolingE ensemble probe across llama-2 response model sizes and test sets across tasks. For each task and model size, a probe is trained on labeled responses of the same task that were generated organically from the 13B model.
Hidden State Type.
Slight variations in grounding saliency—the strength of the signal of hallucination as able to be detected by a probe, measured by probe hallucination classification performance—are also present among hidden states of different types during decoding. As shown in Figure˜3, feed-forward probes reach peak performance at slightly deeper layers than their attention head out-projection probe counterparts and exceed them in single-layer hallucination detection. Training ensemble classifiers on only the hidden states of the same type at every layer, we see how feed-forward hidden states possess a slight but statistically significant greater saliency of model grounding behavior.
@@ 5.3 Results
Training probes in the same manner on the force-decoded hidden states of responses on our dataset using llama-2-7b and llama-2-70b, we find no statistically significant difference (via Kolmogorov–Smirnov) in hallucination detection across the hidden states from base models of varying sizes (Table˜6). These results show that we can still detect a hallucination even when it was originally generated from a different model, and that even the lower-dimensional hidden states of the 7B model still contain enough information to train a successful probe.
Figure 4: Intrinsic and extrinsic hallucination saliency across layers for abstractive summarization as measured by response-level probe recall on hallucination types.
@@
Next, we test for the saliency of different forms of hallucination behavior in model internal states. Here, with annotator-labeled test set responses according to whether they contained intrinsic or extrinsic hallucinations, we evaluate trained probe performance in correctly detecting the presence of hallucinations for instances of each hallucination type. We observe in Figure˜4 that extrinsic hallucination behavior is more salient in the hidden states of shallower layers for both organic and synthetic hallucinations, with saliency noticeably peaking around layers 15 and 16, decreasing in deeper layers. Overall, extrinsic hallucination remained more salient in model hidden states than intrinsic hallucination.
@@
As Table˜7 shows, probes achieve very high F1 in the detection of synthetically created hallucinations across all tasks. However, probe performance is subject to a significant drop when trained and tested on different modalities, with probes generally performing better when trained on organic hallucinations and tested on synthetic ungrounded responses compared to the reverse, with the notable exception being abstractive summarization on the CDM dataset. Prior work has shown that synthetic data generation approaches specifically designed for factuality evaluation Ma et al. (2023) do not align with actual errors made by generation models in their lexical error distributions Goyal and Durrett (2021). Here, our results show this to extend to their hidden state signals as well, which are used by the probe to detect unfaithfulness in the case of forced decoding on human-created content, relative to organic instances of hallucination.
@@
CDM Organic Organic 0.45
Synthetic Organic 0.48
Organic Synthetic 0.38
Synthetic Synthetic 0.93
CF Organic Organic 0.88
Synthetic Organic 0.65
Organic Synthetic 0.83
Synthetic Synthetic 0.94
E2E Organic Organic 0.87
Synthetic Organic 0.69
@@
Synthetic Synthetic 0.98
Table 7: Response-level PoolingE probe performance, as measured by F1-R, in detecting hallucinations on organic or synthetic ungrounded responses. Performance is generally better when the training data resembles the test data.
Task.
To test whether the saliency of grounding behavior persists consistently in similar hidden state locations across different tasks, we train our probe on each specific task and evaluate its performance when applied to the others. As shown in Table˜8 and compared to same-task training, we find a lack of generalization in hallucination detection performance for our probes across different train and test tasks. This may mean that there is not a consistent signal of hallucination across tasks, or that our probes are not distinguishing this signal from its task-or-dataset-specific confounds, and are overfitting to the confounds.
6 Discussion
In this work, we examined how LLM internal representations reveal their hallucinatory behavior, via a case study of several in-context grounded generation tasks. Our analyses point to several nuances in the saliency of this behavior across model layers, hidden state types, model sizes, hallucination types, and contexts. Below, we highlight a few takeaways as well as limitations that naturally suggest directions for future work.

### v2

@@ 3.3 Annotation
We ask annotators to label the minimal hallucinatory spans in each response, if any. These are the smallest text spans that would need modification to transform an ungrounded response into a grounded one. A team of 17 annotators completed a pilot annotation exercise of 50 annotations per task. These pilot annotations were checked and verified by the authors to pinpoint common sources of error and disagreement. This feedback was integrated into a revised set of annotation guidelines, shown in Appendix A, and the full dataset was then re-annotated fully by three annotators such that each example was annotated three times (example annotations are shown in Table 1). Annotations were then reconciled at the token-level into the final gold annotation, such that a token is labeled as a hallucination if at least two out of three annotators included it in a hallucination-span annotation.
Annotation statistics are shown in Table 3. Note that for each dataset, over one-fifth of all annotated responses were annotated as containing at least one hallucinated span (columns H-o and H-s), which we refer to as response-level annotations. Annotators exhibited moderate agreement Landis and Koch (1977), achieving response-level Fleiss’ $\kappa$ scores between 0.39 and 0.55 across the three tasks.
@@ 4 Probing
Probes Alain and Bengio (2017) are tools used to analyze a neural network’s internal representations. Often taking the form of linear classifiers, they are applied to the internal representations of a network and are trained to discriminate between different types of inputs or outputs. Here, we are interested in detecting hallucinatory outputs during an LLM’s generation steps when conditioned on a prompt that requires a response grounded in the prompt.
@@ 4.1 Design
Let $u$ be a prompt as described at the start of § 3. To determine whether a response $x$ to a prompt $u$ contains hallucinations, we will probe the hidden states of the transformer language model (i.e., decoder) as it generates $x$. Note that these hidden states depend on $u$ and on the prefix of $x$ that has been generated so far, and they may be reconstructed from the string pair $(u,x)$ via forced decoding. A single transformer layer actually consists of two sublayers Vaswani et al. (2017): the first transforms a token’s embedding by adding a vector predicted by attention, and the second transforms it again by adding a vector predicted by a feed-forward network. We train a separate probe for each of the $2L$ sublayers, where $L$ is the number of layers.
@@
This probe is a linear classifier using a single hidden state vector (output by some sublayer) as input. Let $x_{i}$ be the $i^{\textrm{th}}$ response token and let $\mathbf{h}_{i}\in\mathbb{R}^{n}$ be the corresponding LLM hidden state at a fixed layer and component. We write $y_{i}\in\{0,1\}$ to indicate whether $x_{i}$ falls within a minimum hallucination span. The linear probe predicts $y_{i}$ by $p(y_{i}=1\mid u,x_{\leq i})=\sigma(\mathbf{w}^{\intercal}\mathbf{h}_{i}+b)$, where $\mathbf{w}\in\mathbb{R}^{n}$ are learned weights, $b\in\mathbb{R}$ is a learned bias, and $\sigma$ is the logistic function.
@@
[…]
@@ 4.2 Training Objectives
To train a probe to predict token-level hallucination labels $y_{i}$, we fit probe parameters to minimize the negative log-likelihood, $\mathcal{L}(\theta)\propto\sum_{(u,x,y)\in\mathcal{D}}\sum_{i=1}^{|x|}-\log p(y_{i}\mid u,x_{\leq i};\theta),$ on the token-level annotations in our dataset $\mathcal{D}$.
@@
We also try training a classifier to predict whether $x$ contains any hallucinations at all; denote this indicator as $y=\bigvee_{i=1}^{|x|}y_{i}\in\{0,1\}$. For this, we simply use an attention-pooling probe over the entire response $x$. To fit the probe parameters, $\theta$, we again minimize negative log-likelihood, $\mathcal{L}(\theta)\propto\sum_{(u,x,y)\in\mathcal{D}}-\log p(y\mid u,x_{\leq|x|};\theta).$
@@ 4.3 Metrics
Span-level Classification.
For span-level classifiers, it is possible to evaluate the F1 score of their per-token predictions; however, this would give more weight to annotated or predicted spans that are longer in length. Here, we are more interested in span-level F1: specifically, whether we predicted all of the annotated hallucinations (recall), and whether all of the predicted spans were annotated as actual hallucinations (precision). As requiring an exact match of spans is too severe—Appendix B shows how our human annotators often disagreed on the exact boundaries of a hallucinatory span—we modify span-level precision and recall metrics to give partial credit, as follows.
Define a span to be a set of tokens, where tokens from different responses are considered distinct. Let $A$ be the set of all annotated spans in the reference corpus, and let $P$ be the set of all predicted spans on the same set of examples. We compute the recall $r$ as the average coverage of each reference span, i.e., the fraction of its tokens that are in some predicted span. Conversely, we compute the precision $p$ as the average coverage of each predicted span. So $r=\frac{1}{|A|}\sum_{s\in A}\frac{\left|s\cap\left(\bigcup_{\hat{s}\in P}\hat{s}\right)\right|}{|s|}$ and $p=\frac{1}{|P|}\sum_{\hat{s}\in P}\frac{\left|\hat{s}\cap\left(\bigcup_{s\in A}s\right)\right|}{|\hat{s}|}$ where $\cup$, $\cap$, and ||$\cdot$|| ⋅ | denote set union, intersection, and cardinality respectively. We define F1-Sp (standing for "span, partial credit") as the harmonic mean of $p$ and $r$. Note that these quantities micro-average over spans, ignoring the boundaries between examples. Thus, in contrast to F1-R, examples with more spans have more influence on the final F1-Sp metric achieved.
@@ 5.1 Probe Hyperparameters and Training
Single-layer hidden state probes, described in § 4, are trained under float32 precision with unregularized log loss. Optimal batch sizes between 20-100 and Adam optimization Kingma and Ba (2014)—with learning rates between 0.1-0.001—and $\beta_{1},\beta_{2}=(0.9,0.999)$ were chosen for each task after a small grid search on hyperparameters. With 70/10/20% train, validation, and test splits, each classifier was trained using early stopping with 10-epoch patience. The ensemble classifier was trained with the same hyperparameters as the single-layer probes.
For CDM, the synthetic training dataset (from BUMP) is smaller than the organic training dataset, so the SYNTH results are at a disadvantage. For CF and E2E, however, we generated a synthetic training dataset that was matched in size.
When reporting results, SL denotes the best single-layer classifier (chosen on the validation set), E denotes the ensembled classifier, and SYNTH indicates that the classifier was trained on synthetic data and tested on organic hallucinations.
@@ 5.2 Response-Level Baselines
We compare the performance of our probes against a set of contemporary methods used in faithfulness evaluation, measuring test set hallucination classification F1. Specifically: (1) the model uncertainty-based metric of length-normalized sequence log-probability Seq-Logprob Guerreiro et al. (2023); (2) off-the-shelf synthetically-trained sentence-level claim classification FactCC Kryscinski et al. (2020); (3) linguistic feature-based dependency arc entailment DAE Goyal and Durrett (2020); (4) QA-based metrics FEQA Durmus et al. (2020); (5) QuestEval Scialom et al. (2021); (6) the NLI-based metric SummaC Laban et al. (2022); ChatGPT classifiers Luo et al. (2023) gpt-3.5-turbo-0125 at zero temperature (7) without and (8) with basic chain-of-thought prompting (with prompts shown in § C.1), as well as (9) the automatic claim-breaking and verification method FActScore Min et al. (2023) with gpt-3.5-turbo-instruct.
Metrics (2)–(6) and (9) were not originally intended to evaluate individual responses, but rather a collection of responses for the sake of system comparisons. To adapt to our setting, however, we threshold each of these continuous metrics to classify the individual responses, setting the threshold to maximize validation set F1, following Laban et al. (2022). For sentence-level optimized baselines such as FactCC, response-level evaluation is performed by logical OR-ing sentence-level baseline predictions (obtained after thresholding the raw scores). In other words, any hallucination in any sentence leads to a "hallucinated" label. Note that, for our abstractive summarization tasks, methods such as FactCC Kryscinski et al. (2020) are explicitly trained in the domain as they use CDM as the initial dataset for training. For certain baselines in CF and E2E, dashes are present in the results due to an inability to directly compare methods in these domains. For example, for DAE, one cannot obtain the dependency arcs from a set of restaurant attributes not expressed in natural language.
5.3 Expert Human Judge
To compare our probe performance to a human performing the equivalent hallucination detection task, we had a another professional annotator judge the test set. This annotator (denoted Expert Human) was distinct from the annotators who produced the gold labels. They were given access to the training data (consisting of all three sets of annotations plus the gold reconciled annotations) and performed a round of pilot annotations with feedback from the authors before annotating the test set.
5.4 Results
@@
OC 0.43 0.79 0.71
Seq-Logprob 0.64 0.81 0.72
FactCC 0.51 - -
DAE 0.53 - -
FEQA 0.44 - -
QuestEval 0.61 - -
SummaCZS 0.68 - -
SummaCConv 0.67 - -
ChatGPT 0.63 0.74 0.74
ChatGPT CoT 0.53 0.51 0.47
FActScore 0.61 0.80 0.72
PoolingSL 0.73 0.93 0.89
PoolingE 0.75 0.94 0.90
PoolingSL-SYNTH 0.65 0.80 0.68
PoolingE-SYNTH 0.66 0.83 0.71
Expert Human 0.55 0.89 0.79
Table 4: Response-level probe and baseline organic hallucination detection performance across tasks as measured via F1 scores (F1-R) achieved on the test set.
For performance comparisons, probe results are statistically significant ($p<0.05$) relative to respective baselines under exact paired-permutation testing Zmigrod et al. (2022) unless otherwise noted.
Response-level Results.
Response-level results are shown in Table 4. On all three tasks, the probes were able to surpass the Expert Human annotator on response-level classification, with the PoolingE probe outperforming the Expert Human by as much as 20 points absolute F1-R on the CDM dataset. The CDM dataset is likely the most challenging for a human to judge as it requires checking the generated statements against the complete (and lengthy) news article. Both PoolingSL and PoolingE probes outperformed all baselines across all datasets. Ensembling generally adds at least one point in absolute F1-R above the best single layer. The probes trained on synthetic hallucinations generally do not perform as well as their organic equivalents, though they are still fairly competitive with the baselines. Seq-Logprob was surprisingly strong, beating both the Expert Human and all but one of the baselines on CDM. This suggests that while model confidence may not be the only feature related to hallucination detection, it is likely an important one.
Span-Level Results.
These results appear in Table 5. The PoolingE probes match or surpass the Expert Human performance on the CDM and E2E datasets while falling three points behind in absolute F1-Sp on the Conv-FEVER dataset. Pooling information across time does help probe performance with the best single-layer pooling probe matching or beating the best single-layer linear probe across all datasets. Interestingly, what the linear probe lacks in time can be made up in depth—LinearE matches PoolingSL on CDM and Conv-FEVER, and beats it by one point absolute F1-Sp on E2E, though this wasn’t statistically significant.
6 Analysis
Layers.
Evaluating the performance of trained single-layer attention-pooling probes (Figures 2 and 4), in model middle layers during decoding. Probe performance shows a consistent and gradual increase from the first layer after token embeddings up until layers 10 to 20. Hallucination information is reliably present at all layers $\geq 5$, although there are local optima around layers 5–10 and again at 40. Ensembling across many layers provides only a small (not stat. sig.) improvement over the best single layer (Table 4), so we do not have evidence that our probes at different layers detect different types or correlates of hallucination. In general, probe performance declines for subsequent layers after their initial peak layer across tasks, though CDM is an exception in that probe performance gradually increases again from layers 30 to 40, peaking there.
Hidden State Type.
Slight variations in grounding saliency—the strength of the signal of hallucination as able to be detected by a probe, measured by probe hallucination classification performance—are also present among hidden states of different types during decoding. As shown in Figure 2, feed-forward probes reach peak performance at slightly deeper layers than their attention head out-projection probe counterparts and exceed them in single-layer hallucination detection. When training ensemble classifiers on only the hidden states of the same type at every layer, feed-forward hidden states possess a slightly greater & statistically significant saliency of model grounding behavior.
@@ 6 Analysis
OC 0.31 0.75 0.13
LinearSL 0.52 0.77 0.54
LinearE 0.54 0.79 0.55
PoolingSL 0.54 0.79 0.54
PoolingE 0.55 0.81 0.56
Expert Human 0.40 0.84 0.56
Table 5: Span-level probe performance, as measured by span-level partial match scores (F1-Sp).
@@
Training probes in the same manner on the force-decoded hidden states of responses on our dataset using llama-2-7b and llama-2-70b, we find no statistically significant difference in hallucination detection across the hidden states from base models of varying sizes. These results show that we can still detect a hallucination even when it was originally generated from a different model, and that even the lower-dimensional hidden states of the 7B model still contain enough information to train a probe.
Figure 2: Grounding behavior saliency as measured by probe response-level F1 (F1-R) on the CDM test set. The $x$-axis corresponds to the layer of the probe. Plots across all tasks are shown in Appendices D and 4. Vertical dashed lines denote the layers that respective probes first surpass 95% of their peak performance.
Figure 3: Intrinsic and extrinsic hallucination saliency across layers for synthetic CDM as measured by response-level MLP probe F1 (F1-R) on hallucination types; additional plots in Appendix D, Figure 5.
@@
Next, we test for the saliency of different forms of hallucination behavior in model internal states. Here, with annotator-labeled test set responses according to whether they contained intrinsic or extrinsic hallucinations,we evaluate trained probe performance in correctly detecting the presence of hallucinations for instances of each hallucination type. Intrinsic hallucinations are rarer and may reflect a misunderstanding of the grounding data. We observe in Figure 3 that extrinsic hallucinations—which are 1.8×1.8$\times$1.8 × more common at the response level (as labeled by annotators) and presumably reflect a failure to find desired information in the grounding data at all—are easier to detect at every layer. Both kinds are generally most salient in probe middle layers, between layers 10 to 20. Overall, extrinsic hallucination remained more salient in hidden states than intrinsic hallucination.
@@
As Table 6 shows, probes achieve very high F1 on the detection of synthetically created hallucinations across all tasks. However, probe performance is subject to a drop when trained and tested on different modalities, with probes generally performing better when trained and tested on the same hallucination modality. The most severe version of this is on CDM, which may be caused by differences between our organic annotation instructions and those of the BUMP dataset Ma et al. (2023). Prior work has shown that synthetic data generation approaches specifically designed for factuality evaluation Ma et al. (2023) do not align with actual errors made by generation models in their lexical error distributions Goyal and Durrett (2021). Here, our results show this to extend to their hidden state signals as well, which are used by the probe to detect unfaithfulness in the case of forced decoding on human-created content, relative to organic instances of hallucination. This is also related to the idea of ecological validity De Vries et al. (2020). Findings from an ecologically valid experiment should hold outside the context of that study. We do not find strong evidence that the detection of synthetic hallucinations generalizes outside that context.
@@
CDM Organic Organic 0.75
Synthetic Organic 0.66
Organic Synthetic 0.17
Synthetic Synthetic 0.86
CF Organic Organic 0.94
Synthetic Organic 0.83
@@
Synthetic Synthetic 0.95
E2E Organic Organic 0.90
Synthetic Organic 0.71
Organic Synthetic 0.94
Synthetic Synthetic 0.99
Table 6: Response-level PoolingE probe performance, as measured by F1-R, in detecting hallucinations on organic or synthetic ungrounded responses.
Lack of Task Transfer.
To test whether the probed features persist in similar hidden state locations across different tasks, we test our probe on each specific task when trained on the others. As shown in Table 8 in Appendix D, we find a lack of generalization in hallucination detection performance for our probes across different training and test tasks. This may only mean that each probe is overfitting to task-or-dataset-specific confounds. If so, training on a mixture of many tasks (more than the two we tried in Table 8) might result in a classifier that identifies the consistent hallucination signal and thus generalizes well to new tasks. However, it is also possible that no consistent hallucination signal exists.
7 Conclusion
We examined how an LLM’s internal representations reveal its hallucinatory behavior, via a case study of multiple in-context grounded generation tasks. In specific domains, we can accurately identify hallucinatory behavior through trained probes. Probe accuracy can depend on model layer, hidden state type, model size, hallucination type, and training-test mismatch. 1TL;DR: Yes, they do.
Limitations
Below, we highlight a few takeaways, broader implications, and limitations that naturally suggest directions for future work.
Hallucinatory Span Detection.
In the context of use cases such as retrieval-augmented generation (RAG), span-level results have the potential to be extremely useful because they allow us to detect where hallucinations may fall. For example, hallucinatory spans could be (1) highlighted in the UI for RAG-based applications, (2) detected as soon as they are generated, triggering the decoder to backtrack and regenerate, or (3) a local hallucination detection could be used for further mitigation of hallucination in a reward-shaping setting for RLHF, in contrast to only having response-level signals available.